OpenStudy (anonymous):

Help Please !! I need someone to show me step by step ! The largest swimming pool in the world is in Casablanca. The rectangular pool has a perimeter of 1110 meters, and its length is 30 meters longer than 6 times its width. How many meters wide is it?

OpenStudy (anonymous):

Hey can you help me ? :) @bigbrain446

OpenStudy (anonymous):

Alrighty then let's get 'er done. We want width right? Let's make width x

OpenStudy (anonymous):

ok so the we would use this formula is p=2w+2l

OpenStudy (anonymous):

yup yup! now plug in your jazz.

OpenStudy (anonymous):

wait a mo though

OpenStudy (anonymous):

we dont know length or width, but we know that length is 6 times the width plus 32

OpenStudy (anonymous):

so what should L be instead?

OpenStudy (anonymous):

(Maybe some variation of W?)

OpenStudy (anonymous):

Am I making sense?

OpenStudy (anonymous):

L= 6w+32

OpenStudy (anonymous):

Right! So that can take the place of L in the earlier equation. Also for simplicity it can be done like this 1/2p = w + (6w+30) since you only need half the perimeter to find the length or width.

OpenStudy (anonymous):

ok so how would i slove that now ?

OpenStudy (anonymous):

W(6w) + w(30) ?

OpenStudy (anonymous):

plug in the perimeter for p in the equation i mentioned 1/2p = w + (6w+30)

OpenStudy (anonymous):

555=w+(6w+30)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Im so confused :(

OpenStudy (anonymous):

Crud did I dig myself a hole? This should be making a whole lot of sense right about now

OpenStudy (anonymous):

wow thats rude !

OpenStudy (anonymous):

No no I mean I think I might have lead you down a wrong path

OpenStudy (anonymous):

I'm really sorry I thought I knew what I was doing. I'll need a minute to figure it out <:/

OpenStudy (anonymous):

need help im pretty sure i got it?

OpenStudy (anonymous):

I started plugging numbers in and your answer is 75. Maybe we can work backwards from there.

OpenStudy (anonymous):

w = 75 L = 6(75)+30 L = 480

OpenStudy (anonymous):

no your doing it wrong

OpenStudy (anonymous):

Show me how please

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

since we know p=2l+2w, and l=6w+30

OpenStudy (anonymous):



OpenStudy (anonymous):

when we add it up we get 14w+60=1110

OpenStudy (anonymous):

so the answer is 75 :)

OpenStudy (anonymous):

we subtract each side by 60 to get 14w=1050

OpenStudy (anonymous):

then 1050 divide by 14

OpenStudy (anonymous):

w=75 :)

OpenStudy (anonymous):

Alright that makes sense. Sorry about the initial flimflam jackie